OpenStudy (anonymous):

A student spins a 15 g metal ball weight at the end of a 0.60 m string with a frequency of 3.0 revolutions per second. What is the angular momentum of the ball?

OpenStudy (anonymous):

The angular momentum is given as I = mvr , where m is the mass of the ball, v is the tangential speed, and r is the radius of the rotation of the ball. The trick is how to find v? Well v is just distance over time. The distance covered is just 3 times the circumference of the circle per second. Hope that helps :)

OpenStudy (shamim):

Angular momentum L=mvr=m*omega*r^2=?

OpenStudy (shamim):

Given m=15g=0.15kg Omega=2*pi*f=? f=3hz r=0.6m

OpenStudy (shamim):

Response plz!!!!!